OpenStudy (anonymous):

The equation of a line is 2x - 2y = 6. What is the slope of the line

OpenStudy (shamil98):

turn the equation into slope-intercept form. y = mx + b

OpenStudy (shamil98):

-2y = -2x + 6 y = x - 3

OpenStudy (anonymous):

so the answer is negative 3 ?

OpenStudy (shamil98):

the slope is 1. y= mx + b 1 is multiplied to x. therefore the slope is 1.

OpenStudy (anonymous):

oh that was my first answer too